# Bits-to-Photon: End-to-End Learned Scalable Point Cloud Compression for Direct Rendering

### v1

@@ Abstract
Keywords:
Scalable Point Cloud Compression Point Cloud Rendering 3D Sparse Convolution
@@ 1 Introduction
Volumetric video streaming is a promising technology with the recent advancement of AR/VR devices. It allows users to view a 3D scene with 6 degrees of freedom (6-DoF) while the scene changes in time. It is expected to revolutionize both the way we communicate and entertain. However, since the 3D scene needs to be captured, transmitted, and rendered in real time, it inevitably incurs high requirements in network bandwidth and computation resources.
@@
To reduce the consumption of bandwidth and alleviate the computational burden of surface reconstruction, point clouds have been recognized as a form of promising 3D representation for volumetric streaming in emerging AR/VR applications [12, 19], given the flexibility in capturing and compression of point clouds. Tremendous progress has been made in the past few years in point cloud compression, including the MPEG Point Cloud Compression (PCC) standards [21] and recent deep learning based methods [26, 31, 14]. These methods are typically developed to compress either the geometry (point coordinates) or color, and have shown great potential in compressing point clouds to a low bit-rate. However, there remains a gap between the reconstruction of point cloud and the rendering of high-quality images on the clients’ display. The color compression methods are mostly optimized to accurately reproduce the point-wise colors, rather than the rendering quality. As a result, the rendering quality from the compressed point clouds using standard rendering tools is sometimes severely compromised with visible holes and color artifacts. Although the state-of-the-art learned point cloud rendering method [4] can improve the rendering quality, it is very slow and cannot support real-time free viewpoint.
In this paper, we propose a point cloud compression scheme that generates a bit stream that can be directly decoded to 3D Gaussians, which can then be rendered efficiently and with high quality using Gaussian splatting [18]. The encoder and decoder are end-to-end optimized towards lower bit-rates and better rendering quality. It allows multiple levels of details at different bit-rate ranges and significantly improves the rendering quality at similar bit-rates, while substantially reducing the decoding and rendering time compared to existing separate point cloud compression and rendering methods. Our method supports real-time color decoding and rendering of high quality point clouds, paving the way towards building interactive 3D streaming applications with free view points.
Since the scalable coding of point cloud geometry has been well studied in the work of SparsePCGC [25], in this paper we focus on coding the color and other rendering related information in a scalable stream. Following the octree coding structure, we first extract features at different levels of the octree. We then compress the features using a hierarchical coding scheme, where the entropy coding at each level is conditioned on the up-sampled features from the previous coarser level. The feature bits at each level are decoded directly to 3D elliptical Gaussians with opacity and color, which can then be efficiently rendered to arbitrary view points using the Gaussian splatting technique [29, 18, 34]. The proposed approach thus allows multiple levels of details at different bit-rate ranges.
@@
Our feature extraction and compression and decompression modules are built upon the sparse 3D convolutional neural network based on the Minkowski Engine [6]. We propose a novel geometry-invariant 3D sparse convolution to address the problem of non-uniform density in the point cloud, which is crucial for color feature extraction and reconstruction. With the help of a differential renderer leveraging the 3D Gaussian representation, we can end-to-end optimize the encoder and decoder jointly towards better rendering quality and lower bit-rates.
@@
We train our model on the training set of the THuman 2.0 dataset [30], which contains point clouds of full-body human in action. We evaluate our method on the testing set of the THuman 2.0 dataset and the 8i Voxelized Full Bodies (8iVFB) dataset [8]. Our method achieves higher perceptual quality (substantially higher in terms of PSNR) in the rendered images at similar bit rates as G-PCC v22 and the state-of-the-art deep learning based point cloud attribute compression method [9]. Meanwhile, our color decoding (using GPU) is substantially faster than both G-PCC (using CPU, cannot benefit from GPU) and the learning-based method (using GPU), at 10s of milliseconds (ms) per frame, compared to seconds. Although the state-of-the art learned geometry decoding method that we leverage [25] is still at 100s of ms, our method for coding the color and rendering-related information paves the way for future real-time volumetric video streaming applications.
@@
  • We propose a novel hierarchical coding scheme for the compressing the color and rendering-related information. It generates a scalable bit-stream and allows multiple levels of details at different bit-rate ranges. Our method (B2P) significantly improves the rendering quality at similar bit-rates compared to both standard and learned methods for point cloud color compression, while substantially reducing the decoding and rendering time.
@@ 2 Related Works
2.1 Point Cloud Compression
@@
With the wide applications of point clouds in autonomous driving, AR/VR, and volumetric video capturing and streaming, point cloud compression has been an active research area in the past few years. The MPEG 3D Graphics Coding group has standardized the video-based (V-PCC) and the geometry-based point cloud codecs (G-PCC) [5, 11]. The V-PCC projects 3D point clouds to 2D images and encodes them by a standard video codec. It is efficient in terms of compression ratio especially for dynamic point clouds (a.k.a. point cloud videos (PCV)), but lacks level-of-detail (LoD) scalability and may introduce 3D artifacts in the reconstructed geometry. The G-PCC, on the other hand, encodes the geometry using an octree structure, with context-adaptive entropy coding. G-PCC stream is scalable, making it easy to adapt to time-varying network bandwidths in streaming applications. Given the coded geometry, it adopts the Region-Adaptive Hierarchical Transform (RAHT) to efficiently encode the color, achieving high rate-distortion performance.
@@
With the advancement in 3D neural network architectures [28, 24, 6], neural network-based methods show great potential in boosting the rate-distortion performance in point cloud compression. One line of works focus on using neural networks to improve conditional entropy modeling on the octree structure for lossless geometry compression [15, 22, 10, 7, 20]. These works are particularly optimized for compression of geometry and are not directly applicable to the color and other rendering related information. Another line of works focus on compressing point clouds sampled from a 3D surface, where the geometry is usually coded in the first place [27, 25, 14] and colors are coded conditioned on the decoded geometry [23, 9, 26, 31]. These methods demonstrate great potential in compressing point clouds to a low bit-rate while maintaining the color and geometry fidelity. However, all these methods compress geometry and color separately, unaware of the rendering quality tradeoff, and sometimes lead to visible holes and color artifacts in the rendered images. In this paper, we propose a novel end-to-end learned method that compresses the color and other rendering information, so that the decoded bit stream along with the geometry bits can be directly decoded to renderable 3D Gaussians, bridging the gap between point cloud compression, reconstruction, and rendering.
2.2 Differential Point Cloud Rendering
Our work is based on the key innovation of differential point cloud rendering [2, 29, 4], which allows us to optimize the bitrate-rendering-distortion tradeoff directly. Specifically, we are inspired by the work of 3D Gaussian splatting [18], which demonstrate that 3D Gaussian clouds are capable of representing a 3D scene with high fidelity and rendering images without holes. Their work focused on the generation of 3D Gaussians from given multi-view images of one scene through per-scene optimization. In this work, we tackle a different problem, where we are given a colored point cloud and aim to compress it to a low bit-rate while maintaining the rendering quality. Our key contributions are developing neural network modules to extract compact features from the point cloud, conduct entropy coding, and decode to 3D Gaussians for high-quality rendering.
@@ 3 Method
3.1 Hierarchical Compression and Rendering
Figure 1: Hierarchical Compression and Rendering framework.
@@ 3.1 Hierarchical Compression and Rendering
We represent the point cloud geometry (i.e. the coordinates of all points) using an octree structure, which has been found to be efficient in representing the sparse structure of point clouds [16]. The octree is constructed by recursively subdividing the 3D space into 8 equal octants. Each node represents the occupancy of the corresponding octant. If one octant is completely empty, the corresponding node is marked unoccupied and not further subdivided. A point cloud where each coordinate is represented with $N$ bits can be represented by an octree with $N$ levels, where each leaf node represents an original point in the point cloud. The octree structure provides a hierarchical representation of the geometry of the point cloud, which naturally facilitates scalable compression and rendering with multiple levels of detail. It also allows us to efficiently find the neighbors of a point in the point cloud, which is crucial for sparse 3D convolutions.
@@
With the octree geometry structure, we compress the point cloud by first losslessly encode the geometry using an existing scalable point cloud geometry compression method, such as SparsePCGC [25] or G-PCC [21], and then encode the color and other rendering related information given the geometry defined by the octree structure. The hierarchical compression and rendering pipeline is illustrated in Fig. 1.
@@
We first extract color features [^1] For simplicity, we call these color features. But they actually carry additional information needed to recover the 3D Gaussians for rendering. from the original point cloud at level $N$. We then generate hierarchical feature representation by repeatedly downsampling the extracted features following the octree structure, until a chosen level $L$, using the average pooling in a $2\times 2\times 2$ neighborhood. We then encode the features at different levels using a hierarchical coding scheme. We start with coding the color features at level $L$ independently, and then code features at successive higher levels conditioned on the upsampled features from the lower level, until level $N$.
@@
The hierarchical coding scheme naturally generates a scalable bit-stream, and allows us to achieve different bit-rates and different levels of details, by transmitting up to different levels $M$ ($L\leq M\leq N$), with $M$ chosen based on the sustainable network throughput between the sender and the receiver or the decoder processing constraint. We have observed that, when the geometry is too sparse (i.e, $M$ is too small), it is very hard for the features to carry enough information for rendering with adequate quality. We found that for a typical 10-bit point cloud ($N=10$), the best rate-quality tradeoff is achieved by transmitting starting from level $L=7$ up to level $M\in\{8,9\}$ and rendering at level $M$. Although it is possible to transmit all way to level $10$, the quality improvement from level $9$ to level $10$ does not justify the additional bits needed.
@@
We extract features at level $N$ using a multi-layer Minkowski Convolution Network [6]. The original Minkowski convolution is defined as,
@@
where $K$ is the kernel size and $\mathcal{N}(u,K,P^{\text{in}})$ is the set of occupied input points that are within the kernel size of the output point $u$. $P^{\text{in}}$ and $P^{\text{out}}$ denote the set of occupied points. We observe that this definition of convolution is not geometry invariant, as it tends to produce higher magnitude where the neighborhood around $u$ is denser. To address this problem, we propose a geometry-invariant 3D sparse convolution, which is crucial for color feature extraction and reconstruction. The geometry-invariant 3D sparse convolution is defined as,
@@
$\mathbf{x}_{u}=\frac{1}{\sqrt{\sum_{u+i\in\mathcal{N}(u,K,C^{\text{in}})}W_{i}^{2}}}\sum_{u+i\in\mathcal{N}(u,K,C^{\text{in}})}W_{i}\mathbf{x}_{u+i}\text{ for }u\in C^{out},$ (2)
@@
which normalizes the output response by the norm of active kernel weights. This normalization ensures that the convolution response is consistent across different point densities in the local area.
We detail the conditional entropy coding in Sec. 3.2 and the 3D Gaussian decoding in Sec. 3.3. Please refer to the supplementary material for more details on the network architectures we used to extract features and generate the 3D Gaussian clouds.
3.2 Feature Squeezing and Conditional Entropy Coding
Figure 2: Predictive entropy coding for level $n$, conditioned on the reconstructed feature at a coarser level $n-1$.
@@ 3.2 Feature Squeezing and Conditional Entropy Coding
We compress the features at each level using feature squeezing and conditional entropy coding, shown in Fig. 2. Suppose we are coding input features at level $n$, given reconstructed features already decoded at level $n-1$. We first apply a feature squeezing module to reduce the number of input feature channels from 64 to 8. The motivation of feature squeezing is that the features at level $n$ can usually be predicted well from features at level $n-1$ and we only need to encode the residual information in level $n$, which is usually sparser and requires fewer channels to represent. The squeezing module learns to generate the residual in a general sense given the context, and prepare the residual features for quantization and entropy coding. We use the Minkowski Inception ResNet architecture [27] with the proposed geometry-invariant 3D convolution for the squeezing module.
@@
To encode the squeezed features, we first quantize the feature components to integers, and use a conditional arithmetic coder to encode the quantized features to binary bit-streams using the up-sampled features from the reconstructed features at level $n-1$ as the conditioning context. During training, the quantization is simulated using additive uniform noise $U(-0.5,0.5)$. Following the entropy coding method of [3], we assume the unquantized feature values follow a Gaussian distribution with spatially varying mean and variance, so that the probability of each quantized value $x_{u}$ at point $u$ is calculated as,
$p(X=x_{u})=\Phi(x_{u}+0.5;\mu_{u},\sigma_{u})-\Phi(x_{u}-0.5;\mu_{u},\sigma_{u}),$ (3)
@@
where $\Phi$ is the Gaussian cumulative distribution function (CDF) and $\mu_{u},\sigma_{u}$ are estimated from the context features. We assume the number of bits needed to code $X=x_{u}$ equals the negative log likelihood, as,
$R(x_{u})=-\log_{2}p(X=x_{u}).$ (4)
Hence, we define the bit-rate loss at level $n$ as the sum of the bits for all points at this level $P_{n}$:
$\mathcal{L}_{R}^{n}=\sum_{u\in P_{n}}-\log_{2}p(x_{u}).$ (5)
@@
During inference, the arithmetic coder will utilize the estimated probability $p(X=x_{u})$ from the context features to encode the value $X=x_{u}$, with a code length very close to (4) .
@@
After entropy coding, the quantized features are concatenated with the context features along the channel dimension, and go through the feature reconstruction module. This module produces the reconstructed features at level $n$, which serve two purposes: (1) It can be used as the context to encode the next level $n+1$; (2) If we need to render at level $n$, we can generate the 3D Gaussian cloud directly from the reconstructed features. By encoding and decoding recursively across the scale using the proposed scheme, we achieve level-of-detail scalability, while reducing the total bit rate up to level $N$.
3.3 From Features to 3D Gaussians
Figure 3: Generative rendering.
@@ 3.3 From Features to 3D Gaussians
The reconstructed features we obtain at each level are feature vectors attached to some grid points in the voxelized space. We convert the feature vectors at the highest level $M$ that is transmitted and received to renderable 3D Gaussians. We use the differential Gaussian splatting renderer proposed in [18]. Each 3D Gaussian is parameterized by a center coordinate (a.k.a. the 3D means) $(\mu_{x},\mu_{y},\mu_{z})$, a covariance matrix $\Sigma$, an opacity value $o$, and a color triplet $c$. Since the features are extracted from the original point cloud that does not have anisotropic color, we simplify the original parameterization by using color $c\in[0,1]^{3}$ instead of the spherical harmonics. Each 3D Gaussian is then rendered to the screen space using the splatting technique, and the rendered colors by overlapping Gaussians are combined using alpha blending. A pixel in the screen space is shaded as color $C$, where,
$C=\sum_{k=1}^{K}T_{k}\alpha_{k}c_{k},$ (6)
where $\alpha_{k}$ is the opacity of the $k$-th splat, and $T_{k}=\prod_{j=1}^{k-1}(1-\alpha_{j})$. The opacity $\alpha_{k}$ is calculated by projecting the 3D Gaussian onto the screen space and evaluate the Gaussian density at the projected pixel, weighted by the estimated opacity value $o_{k}$,
$\alpha_{k}=o_{k}\cdot e^{-\frac{1}{2}\mathbf{d}_{k}^{T}\Sigma_{k(s)}^{-1}\mathbf{d}_{k}},$ (7)
where $\Sigma_{k(s)}$ denotes the screen space covariance matrix and $\mathbf{d}_{k}$ is the distance vector between the pixel and the Gaussian center in the screen space. The screen space covariance matrix $\Sigma_{k(s)}$ can be approximated as,
$\Sigma_{k(s)}=\mathbf{J}\mathbf{R}_{c}\Sigma_{k}\mathbf{R}_{c}^{T}\mathbf{J}^{T},$ (8)
where $\mathbf{J}$ is the Jacobian matrix of the projection from the 3D space to the screen space, and $\mathbf{R}_{c}$ is the rotation matrix of the camera. We follow the parameterization of the 3D Gaussian as in [18], where the covariance matrix is parameterized as $\Sigma=\mathbf{R}^{T}\mathbf{S}^{T}\mathbf{S}\mathbf{R}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}=\text{diag}(\sigma_{x},\sigma_{y},\sigma_{z})$ is a diagonal matrix. This parameterization guarantees $\Sigma$ to be positive semi-definite. The rotation matrix $\mathbf{R}$ is calculated from a quaternion $\mathbf{q}=(q_{w},q_{x},q_{y},q_{z})$. The conversion from the reconstructed features to 3D Gaussians is in effect generating the following 14 parameters, $\mu_{x},\mu_{y},\mu_{z},\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},o,c$, for all non-empty points.
@@
A simple approach for this conversion is just to use the grid coordinates of the points as the 3D means, and predict the remaining parameters from the features. However, this simple approach has two limitations. First, the number of grid points decrease exponentially with the level of the octree. At a coarse level there might not be enough Gaussians to render the fine-grained texture of the 3D scene. Second, at a coarse level the grid coordinates may not be the correct center positions due to octree-based downsampling. Therefore, we propose a generative rendering scheme, shown in Fig. 3. The key idea is to generate more points from the original grid points, while allowing the module learning to disable falsely generated points by setting the opacity $o\to 0$. For each grid point with a feature vector, we generate features at 8 sub-points using the generative transpose 3D convolution. The features at all subpoints are then converted to predicted Gaussian parameters, including the coordinate offsets $(\delta_{x},\delta_{y},\delta_{z})$, and other parameters $\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},o,c$. The coordinate offsets are added to the grid point coordinates to produce the 3D means $(\mu_{x},\mu_{y},\mu_{z})$. Using the information carried by the feature, the module learns to enable or disable (by setting opacity value close to 0) the sub-points and move them to the correct positions through the predicted offsets to represent the underlying textured smooth surface.

### v2

@@ 1 Introduction
Volumetric video streaming enables users wearing AR/VR devices to view a 3D scene with 6 degrees of freedom (6-DoF) while the scene changes in time. It is expected to revolutionize both the way we communicate and entertain. Despite the great potential, the high bandwidth requirement and the computational complexity of volumetric video streaming have been major roadblocks to its wide adoption. The 3D scene needs to be transmitted and rendered in real time, which inevitably incurs high requirements in network bandwidth and computation resources.
@@
To reduce the consumption of bandwidth and alleviate the computational burden of surface reconstruction, point clouds have been recognized as a promising 3D representation for volumetric streaming in emerging AR/VR applications (Han, Liu, and Qian 2020; Lee et al. 2020), given the flexibility in capturing (Reimat et al. 2021) and compressing (Nakagami et al. 2023) point clouds. Tremendous progress has been made in the past few years in point cloud compression, including the MPEG Point Cloud Compression (PCC) standards (Nakagami et al. 2023) and recent deep learning based methods (Wang and Ma 2022; Zhang et al. 2023a; He et al. 2022). These methods are typically developed to compress either the geometry (point coordinates) or color, and have demonstrated gains over the MPEG PCC standards. However, there remains a gap between the reconstruction of point cloud and the rendering of high-quality images on the clients’ display. The geometry and color compression methods are mostly optimized to accurately reproduce the coordinates and colors of points, rather than the rendering quality. As a result, the rendering quality from the compressed point clouds using standard rendering tools is sometimes severely compromised with visible holes and color artifacts. Although the state-of-the-art learned point cloud rendering method (Chang et al. 2023) can improve the rendering quality, it is very slow and cannot support real-time free viewpoint.
To address the limitations in existing approaches for point cloud compression and rendering, and inspired by the recent work of 3D Gaussian splatting (Kerbl et al. 2023), we develop a novel point cloud compression scheme. Our key innovation in this paper is to design a point cloud compression scheme that compress the point cloud to a compact bit-stream that can be directly decoded to renderable 3D Gaussians, bridging the gap between point cloud compression, reconstruction, and rendering. Because the state-of-the-art scalable geometry coding approach in  (Wang et al. 2022) already achieves high coding performance for geometry, we focus on the scalable compression of the color as well as other information needed for decoding the Gaussian representation. Our method supports real-time decoding and rendering of high quality color point clouds, paving the way towards building interactive 3D streaming applications with free view points.
@@
We achieve the goal of efficiency and high quality by introducing multiple techniques, including a hierarchical compression scheme and a novel geometry-invariant 3D sparse convolution. The hierarchical design allows multiple levels of details at different bit-rate ranges by extracting and coding the Gaussian features following the octree structure, where the feature transformation and entropy coding at a current level are conditioned on the up-sampled features from the previous coarser level. Our feature extraction, compression and decompression modules are built upon the sparse 3D convolutional neural network based on the Minkowski Engine (Choy, Gwak, and Savarese 2019). To overcome the challenge of irregularity with point cloud data, we propose a novel geometry-invariant 3D sparse convolution, which is crucial for color feature extraction and reconstruction. With the help of a fast and differentiable renderer leveraging the 3D Gaussian representation, we can end-to-end optimize the encoder and decoder jointly towards better rendering quality and lower bit-rates.
@@
We train our models using the training set of the THuman 2.0 dataset (Yu et al. 2021) and evaluate them on the testing set of the THuman 2.0 dataset and the 8i Voxelized Full Bodies (8iVFB) dataset (d’Eon et al. 2017). Our method achieves higher perceptual quality (both in terms of PSNR and LPIPS) in the rendered images at similar bit rates as G-PCC v22 and the state-of-the-art deep learning based point cloud methods using the geometry compression model in (Wang et al. 2022) and attribute compression method in (Fang et al. 2022) and (Zhang et al. 2023b). The contributions of our work are summarized as follows:
@@
  • We propose a novel multi-resolution coding framework for compressing the color and rendering-related information. The features at a current resolution are squeezed and entropy-coded and reconstructed conditioned on the features from the lower resolution to maximally exploit the redundancy across resolutions. It generates a scalable bit-stream that provides a higher level of details as the bit rate increases.
  • Our method significantly improves the rendering quality at similar bit-rates compared to both standard and learned methods for point cloud color compression, while substantially reducing the decoding and rendering time.
@@ 2 Related Works
Point Cloud Compression
@@
With the wide applications of point clouds in autonomous driving, AR/VR, and volumetric video streaming, point cloud compression has been an active research area in the past few years. The MPEG 3D Graphics Coding group has standardized the video-based (V-PCC) and the geometry-based point cloud codecs (G-PCC) (Chen et al. 2023; Graziosi et al. 2020). The V-PCC projects 3D point clouds to 2D images and encodes them by a standard video codec. It is efficient in terms of compression ratio especially for point cloud videos, but lacks level-of-detail (LoD) scalability and may introduce 3D artifacts in the reconstructed geometry. The G-PCC, on the other hand, encodes the geometry using an octree structure, with context-adaptive entropy coding. G-PCC stream is scalable, making it easy to adapt to time-varying network bandwidths in streaming applications. Given the coded geometry, it adopts the Region-Adaptive Hierarchical Transform (RAHT) to efficiently encode the color into a scalable stream.
@@
With the advancement in neural network architectures for processing point clouds (Wang et al. 2019; Thomas et al. 2019; Choy, Gwak, and Savarese 2019), neural network-based methods have shown great potential in boosting the rate-distortion performance in point cloud compression. Existing methods either focus on the compression of only point cloud geometry (Huang et al. 2020; Que, Lu, and Xu 2021; Fu et al. 2022; Cui et al. 2023; Mao, Hu, and Wang 2022; Wang et al. 2021, 2022; He et al. 2022) or compression of color given usually losslessly coded geometry (Sheng et al. 2021; Fang et al. 2022; Wang and Ma 2022; Zhang et al. 2023a, b). These methods demonstrate great potential in compressing point clouds to a low bit-rate while maintaining the color and geometry fidelity. However, all these methods compress geometry and color separately, unaware of the rate vs. rendering-quality trade-off, and may lead to blurred details with standard rendering tools. In this paper, we propose a novel end-to-end learned framework that compresses the color and other rendering information, so that the decoded bit stream along with the geometry bits can be directly decoded to renderable 3D Gaussians, leading to higher rate-distortion performance and efficiency.
3D Gaussian Representation for Point Cloud Rendering
Our work is inspired by the work in (Kerbl et al. 2023), which demonstrates that 3D Gaussian clouds are capable of representing a 3D scene with high fidelity and rendering images without holes. Based on this, the recent work of P2ENet (Hu et al. 2024) demonstrates that point cloud can be rendered in high quality and with high speed when converted to 3D Gaussians using a pre-trained neural network without per-scene optimization. This technique has been also used in 3D scene generation tasks (Xu et al. 2024). Since no per-scene optimization process is needed, this technique potentially supports real-time point cloud streaming applications. However, bandwidth limit is still a major roadblock for these applications. Hence in this work, we investigate how to leverage the Gaussian splatting renderer to optimize a point cloud compression network to achieve good trade-off between rate and rendering-quality. We develop neural network modules to extract compact features from the point cloud, conduct entropy coding, and decode the features to 3D Gaussians for high-quality and fast rendering.
@@ 3 Method
Overview of the Proposed Framework
Figure 1: The proposed multi-resolution compression and rendering framework. The source point cloud is encoded into a scalable bit-stream, and delievered up to a resolution based on the sustainable network throughput. The client decodes the received bit-stream to 3D Gaussians and renders the point cloud according to the client current view point.
@@
We represent the point cloud geometry (i.e. the coordinates of all points) using the octree structure, which has been found to be efficient in representing the sparse structure of point clouds (Huang et al. 2006). The octree is constructed by recursively subdividing the 3D space into 8 equal octants. Each node represents the occupancy of the corresponding octant. If one octant is completely empty, the corresponding node is marked unoccupied and not further subdivided. A point cloud where each coordinate is represented with $N$ bits can be represented by an octree with $N$ levels, where each leaf node represents an original point in the point cloud. The octree structure provides a multi-resolution representation of the geometry of the point cloud, which naturally facilitates scalable compression and rendering with multiple levels of detail. It also allows us to efficiently find the neighbors of a point in the point cloud, which is crucial for sparse 3D convolutions.
@@
With the octree structure, we compress the point cloud by first losslessly encode the geometry using an existing scalable point cloud geometry compression method, such as SparsePCGC (Wang et al. 2022) or G-PCC (Nakagami et al. 2023), and then encode the color and other rendering related information given the geometry defined by the octree structure. The proposed compression and rendering pipeline is illustrated in Fig. 1.
@@
We first extract color features  [^1] For simplicity, we call them color features. But they actually carry additional information to recover the 3D Gaussians for rendering. at all leaf nodes at level $N$ of the octree representation of a point cloud. We then generate the hierarchical feature representation by repeatedly downsampling the extracted features following the octree structure, until a chosen level $L$, using the average pooling in a $2\times 2\times 2$ neighborhood. We then encode the features at different levels using a hierarchical coding scheme. We start with coding the color features at level $L$ independently, and then code features at successive higher levels conditioned on the upsampled features from the lower level, until level $N$, using the conditional feature transform and entropy coding modules shown in Fig. 1.
@@
The hierarchical coding scheme naturally generates a scalable bit-stream, and allows us to achieve different bit-rates and different levels of details, by transmitting up to different levels $M$ ($L\leq M\leq N$), with $M$ chosen based on the sustainable network throughput between the sender and the receiver or the decoder processing constraint. At the receiver, the decoded features at level $M$ are then converted to Gaussian parameters using the Gaussian Generation module in Fig. 1.
The individual modules are explained in more detail below. The network architectures used for different modules are given in the supplementary material.
Full-Resolution feature Extraction
@@
We extract features at level $N$ using a multi-layer Minkowski Convolution Network (Choy, Gwak, and Savarese 2019). The original Minkowski convolution is defined as,
@@
where $\mathcal{N}(u,K,P^{\text{in}})$ is the set of occupied input points that are within the $K\times K\times K$ neighborhood of the output point $u$. $P^{\text{in}}$ and $P^{\text{out}}$ denote the set of occupied points. We observe that this definition of convolution is not geometry invariant, as it tends to produce higher magnitude where the neighborhood around $u$ is denser. To address this problem, we propose a geometry-invariant 3D sparse convolution, which is crucial for color feature extraction and reconstruction. The geometry-invariant 3D sparse convolution is defined as,
@@
$\begin{split}&\mathbf{x}_{u}=\left(\sum_{u+i\in\mathcal{N}(u,K,C^{\text{in}})}W_{i}^{2}\right)^{-\frac{1}{2}}\sum_{u+i\in\mathcal{N}(u,K,C^{\text{in}})}W_{i}\mathbf{x}_{u+i}\\ &\text{ for }u\in P^{out},\end{split}$ (2)
@@
which normalizes the output response by the $L_{2}$ norm of active kernel weights. This normalization ensures that the convolution response is consistent across different point densities in the local area.
Conditional Transform and Entropy Coding
Figure 2: Illustration of key components of the proposed method.
@@
We compress the features at each level using feature squeezing and conditional entropy coding, shown in Fig. 2. Suppose we are coding input features at level $n$, given decoded features at level $n-1$. We first apply a feature squeezing module to reduce the number of output feature channels from 64 to 8, by processing both the 64 input channels at level $n$ and the 64 context channels obtained by upsampling the features from level $n-1$. The motivation of feature squeezing is that the features at level $n$ can usually be predicted well from features at level $n-1$ and we only need to encode the residual information in level $n$, which is usually sparser and requires fewer channels to represent. The squeezing module learns to generate the residual in a general sense given the context, and prepare the residual features for quantization and entropy coding. We use the Minkowski Inception ResNet architecture (Wang et al. 2021) with the proposed geometry-invariant 3D convolution for the squeezing module.
@@
To encode the squeezed features, we first quantize the feature components to integers, and use a conditional arithmetic coder to encode the quantized features to binary bit-streams using the up-sampled features from level $n-1$ as the conditioning context. During training, the quantization is simulated using additive uniform noise $U(-0.5,0.5)$. Following the entropy coding method of (Ballé et al. 2018), we assume the all feature values are independent and each follow a Gaussian distribution with spatially varying mean and variance. Let $x_{u}^{d}$ represent the $d$-th component of the $D$ dimensional feature $x_{u}$ at point $u$. The probability of $x_{u}^{d}$ can be calculated as
$p(x_{u}^{d})=\Phi(x_{u}^{d}+0.5;\mu_{u}^{d},\sigma_{u}^{d})-\Phi(x_{u}^{d}-0.5;\mu_{u}^{d},\sigma_{u}^{d}),$ (3)
@@
where $\Phi$ is the Gaussian cumulative distribution function (CDF) and $\mu_{u}^{d},\sigma_{u}^{d}$ are estimated from the context features. We assume the number of bits needed to code $x_{u}^{d}$ equals the negative log likelihood, i.e.,
$R(x_{u}^{d})=-\log_{2}p(x_{u}^{d}).$ (4)
The bit-rate loss at level $n$ is defined as the sum of the bits for all points at this level $P_{n}$:
$\mathcal{L}_{R}^{n}=-\sum_{u\in P_{n}}\sum_{d=1}^{D}\log_{2}p(x_{u}^{d}).$ (5)
@@
During inference, the arithmetic coder will utilize the estimated probability $p(x_{u}^{d})$ from the context features to encode $x_{u}^{d}$, with a code length very close to (4) .
@@
After entropy decoding, the quantized features are concatenated with the context features along the channel dimension, and go through the feature reconstruction module. This module produces the decoded features at level $n$, which serve two purposes: (1) They can be used as the context to encode the next level $n+1$; (2) If we need to render at level $n$, we can generate the 3D Gaussian cloud directly from the reconstructed features. By encoding and decoding recursively across the scale using the proposed scheme, we achieve level-of-detail scalability, while reducing the total bit rate up to level $N$.
Predictive 3D Gaussian Generation
@@
The reconstructed features we obtain at each level are feature vectors attached to occupied grid points in the voxelized space. We convert the feature vectors at the highest level $M$ that is transmitted and received to renderable 3D Gaussians. Each 3D Gaussian is parameterized by a center coordinate (a.k.a. the 3D means) $(\mu_{x},\mu_{y},\mu_{z})$, a covariance matrix $\Sigma$, an opacity value $o$, and a color triplet $c\in[0,1]^{3}$. The covariance matrix is parameterized as $\Sigma=\mathbf{R}^{T}\mathbf{S}^{T}\mathbf{S}\mathbf{R}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}=\text{diag}(\sigma_{x},\sigma_{y},\sigma_{z})$ is a diagonal matrix. The rotation matrix $\mathbf{R}$ is calculated from a quaternion $\mathbf{q}=(q_{w},q_{x},q_{y},q_{z})$. Since the features are extracted from the original point cloud that does not have anisotropic color, we simplify the original parameterization in (Kerbl et al. 2023) by using a constant color $c\in[0,1]^{3}$ for each Gaussian, instead of multiple coefficients of the spherical harmonics. The conversion from the reconstructed features to 3D Gaussians is in effect generating the following 14 parameters, $\mu_{x},\mu_{y},\mu_{z},\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},o,c$, for all non-empty points. Each 3D Gaussian is rendered to the screen space using the differentiable Gaussian splatting renderer proposed in (Kerbl et al. 2023), and the pixels covered by overlapping Gaussian splats are given a color using alpha blending following (Hu et al. 2024).
@@
A simple approach for this conversion is just to use the grid coordinates of the points as the 3D means, and predict the remaining parameters from the features. However, this simple approach has two limitations. First, the number of grid points decrease exponentially with the level of the octree. At a coarse level there might not be enough Gaussians to render the fine-grained texture of the 3D scene. Second, at a coarse level the grid coordinates may not be the correct center positions due to octree-based downsampling. Therefore, we propose a Gaussian Generation module, shown in Fig. 2. The key idea is to generate more points from the original grid points, while allowing the module learning to disable falsely generated points by setting the opacity $o\to 0$. For each grid point with a feature vector, we generate features at $k=8$ sub-points using the generative transpose 3D convolution shown in Fig. 2. The features at all sub-points are then converted to predicted Gaussian parameters, including the coordinate offsets $(\delta_{x},\delta_{y},\delta_{z})$, and other parameters $\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},o,c$. The coordinate offsets are added to the grid point coordinates to produce the 3D means $(\mu_{x},\mu_{y},\mu_{z})$. Using the information carried by the feature, the module learns to enable or disable (by setting opacity value $o$ close to 0) the sub-points and move them to the correct positions through the predicted offsets to represent the underlying textured smooth surface. In Fig. 2, $\delta$ indicates the number of Gaussians with $o$ close to 0.